OpenStudy (enricoanders):

Find the equation (in terms of x) of the line through the points (-1,-1) and (4,2)

OpenStudy (enricoanders):

@ageta

OpenStudy (ageta):

i will be right back

OpenStudy (ageta):

m =(y2-y2)/(x2-x1) let P1(-1,-1) & P2(4,2).

563blackghost (563blackghost):

Would this need to be in slope-intercept or point-slope form?

OpenStudy (ageta):

no

563blackghost (563blackghost):

As @ageta showed we would first find the slope of the two points.... \(\huge{\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}=slope\) So we input the first and second points.... \(\huge{\frac{2-(-1)}{4-(-1)}}=slope\)

OpenStudy (enricoanders):

thats all the question says, i have no idea

OpenStudy (ageta):

@563blackghost help him im busy

OpenStudy (ageta):

no that is not correct we are not using slope formula

OpenStudy (enricoanders):

the slope is 3/5

OpenStudy (ageta):

bye guys

563blackghost (563blackghost):

From what I saw you had used the slope fomula to find the slope @ageta

563blackghost (563blackghost):

Correct :) So we now know the slope....since we are given the first two points and not knowing the y-intercept lets set this in point-slope form.... \(\huge{y-y_{1}=m(x-x_{1})}\) m is the slope....x1 and y1 is the first point which would be `(-1,-1)`....Input these ;)

OpenStudy (enricoanders):

I did that already and tried submitting as y=mx+b , its wrong

563blackghost (563blackghost):

So you submitted in slope-intercept form?

OpenStudy (enricoanders):

yes

563blackghost (563blackghost):

If its not slope-intercept then it would be point-slope form which is what I had shown above.

OpenStudy (enricoanders):

Okay, let me try

OpenStudy (enricoanders):

where you're supposed to put the answer it has y= and then a place for u to put the answer

OpenStudy (enricoanders):

so like y=_________ so no way to put point slope

563blackghost (563blackghost):

Then it would be in slope-intercept form....what was the answer you got before?

OpenStudy (enricoanders):

with the points (-1,-1)(5,-4) y=-1/3x+2/3

OpenStudy (enricoanders):

the points change everytime i try an answer

563blackghost (563blackghost):

We know of the slope which is \(\Large{\frac{3}{5}}\).....but we need to place in slope-intercept form so we need to find the y-intercept....So lets input the slope into the format... \(\huge{y=\frac{3}{5}x+b}\) Now we need to find b...in order to do that we would input one point where x and y is....lets use `(-1,-1)`..... \(\huge{-1=\frac{3}{5}(-1)+b}\) Simplify.... What would b equal?

OpenStudy (enricoanders):

2/5=b

563blackghost (563blackghost):

Close. It would actually be negative instead of a positive. So now we know of the slope and y-intercept so we input it into the slope-intercept format.... \(\huge{y=\frac{3}{5}x-\frac{2}{5}}\) This is the equation.

OpenStudy (enricoanders):

okay let me try haha

563blackghost (563blackghost):

Was it successful?

OpenStudy (enricoanders):

Nope... ://

563blackghost (563blackghost):

hmmmm how did you type it?

OpenStudy (enricoanders):

OMGGG IT WORKED WITH ANOTHER POINT!!! THANK YOU THANK YOU!!!!!

563blackghost (563blackghost):

You welcome! :D Glad it worked!

OpenStudy (enricoanders):

:D